OpenStudy (anonymous):

Algebra 1 - Inverse Variation Suppose a runner takes 45 min to run a route at 8 mi/h at the beginning of training season. By the end of training season, she can run the same route in 38 min. What is her speed at the end of the training season. Someone please HELP me solve this. Don't just tell me the answer.

OpenStudy (anonymous):

Haha, I just asked this question.

OpenStudy (anonymous):

OK, convert to hours and multiply by the initial speed, then you have the distance, then divide that by the later time for her final speed.

OpenStudy (anonymous):

You need to find the distance first. The formula for distance is d = r * t. Where r is the rate (8 mi/hr) and t is the time, 45 minutes. solve for d. Once you find distance (d). Use it for the equation to find rate. The rate equation would be r - d/t. Do you want me to work out the problem with the answer?

OpenStudy (anonymous):

Okay, thanks. I think I got it now. (:

OpenStudy (anonymous):

Btw you need to convert 45 minutes to hours. 45 mins x 1 hour / 60 mins = .75 hours. Use that for your t. No problem

OpenStudy (anonymous):

r t = d, r = d/t\[\frac{\left(8\frac{45 }{60}\right)}{\frac{38}{60}}\text{=}\frac{180}{19}=9.47 \text{ mph}\]